How do i graph this equation by its intercepts.
8x-4y=32

Answer by wuwei96815(245)

This is a linear equation so the graph will be a straight line. To draw a straight line you only need two points.

To find the x-intercept let y = 0
8x-4(0)=32
8x=32
x=32/8=4
Therefore, the x-intercept is the point (4,0)

To find the y-intercept let x = 0
8(0)-4y=32
-4y=32
y=-8
Therefore, the y-intercept is (0,-8)

Plot these two points on a graph and draw a line through them and your job is done.